OpenStudy (anonymous):

how do i simplify the expression √16r^4s^5

OpenStudy (anonymous):

4r^2s^2 radical(s)

OpenStudy (anonymous):

yes but how do i do that?

OpenStudy (anonymous):

\[\sqrt{16r^4s^5}\] \[4r^2\sqrt{s^5}\] \[4r^2 s^{5/2}\]

myininaya (myininaya):

\[\sqrt{16r^4s^5}=\sqrt{16}\sqrt{r^4}\sqrt{s^5}=4r^2\sqrt{s^4s}=4r^2\sqrt{s^4}\sqrt{s}=4r^2s^2\sqrt{s}\]

OpenStudy (anonymous):

1) the square root of 16 is 4 2) the square root of r^4 is r^2 because 2 goes in to 2 twice 3) the square root of s^5 is \[s^2\sqrt{s}\] because 2 goes in to 5 two times, with a remainder of 1

OpenStudy (anonymous):

explanation is written out by myininaya but i am certain she does it in her head the way i wrote it

myininaya (myininaya):

lol no i do it in the way of the myininaya method

OpenStudy (anonymous):

thank you :)